OpenStudy (ahsome):

Find a and b if (a+bi)^2=12+16i. I don't know what to do

OpenStudy (xapproachesinfinity):

well first try to expand that left hand side expression!

OpenStudy (ahsome):

I got a^2+2abi-b^2=12+16i

OpenStudy (xapproachesinfinity):

Good now put some term together \(a^2-b^2+2abi=12+16i\) then \((a^2-b^2)+2abi=12+16i\) Now apply the rule of equality of two complex numbers \(a+bi=c+di\) then \(a=c ~~~and~~~ b=d\)

OpenStudy (ahsome):

So a^2-b^2 = 12, 2ab=16

OpenStudy (xapproachesinfinity):

good those are you two equations solve for a and b

OpenStudy (ahsome):

THANK YOU :D

OpenStudy (xapproachesinfinity):

welcome